# Do assumptions about the central density of subhaloes affect dark matter annihilation and lensing calculations?

### v2

@@ 3 Evolution of the concentration parameter
Density profiles can be characterized by a concentration parameter, which reflects the distribution of mass within the system. For isolated haloes, the concentration parameter $c$ is traditionally defined in terms of the radius at which the logarithmic density profile has a slope of $-2$, i.e., $c=r_{\rm vir}/r_{-2}$ [63], where $r_{\rm vir}$ is the virial radius. As discussed in [52, 29], this definition can be problematic for general cases, as it does not capture deviations from an NFW profile. Several other definitions of concentration exist, but the relationship between them is profile-dependent. Thus we will restrict ourselves to the traditional definition, taking the ratio of $r_{\rm vir}$ and $r_{-2}$.
@@
In cosmological simulations, the virial radius is normally defined in terms of a (redshift-dependent) mean enclosed density. In idealized merger studies, the definition of the virial radius is somewhat arbitrary, but as in Section 2.3, we assume the satellite haloes begin merging with a concentration of $c=10$. This corresponds to an over-density of $\bar{\rho}_{\rm vir}=0.2387\,\rho_{\rm unit}$. As the satellite is stripped, we solve for $r_{\rm vir}$ as the radius where the $\bar{\rho}(r_{\rm vir})=\bar{\rho}_{\rm vir}$. The scale radius, $r_{-2}$, is calculated directly by numerically differentiating and then solving ${\rm d}\log_{10}\rho/{\rm d}\log_{10}r=-2$.
@@
As we will show below, this first definition of concentration generally increases as mass is lost since the scale radius shrinks faster than $r_{\rm vir}$. While this definition is useful for tracking how the bound remnant of a subhalo increases in mean density as it loses its lower-density outer material, $r_{\rm vir}$ has no clear physical significance once a halo has become a subhalo since its density no longer tracks the original cosmological mean value. Thus, in addition to the classical concentration, $c=r_{\rm vir}/r_{-2}$, for stripped systems we will consider an “effective" concentration parameter, $c_{t}=r_{t}/r_{-2}$. As argued in [4, 68], this definition may be more natural as it better reflects the higher mean density of subhaloes and the lower concentration the subhalo had when was originally accreted. In general, we define the tidal radius $r_{t}$ to be the radius at which the density profile is zero. With this definition, the H03 profile does not have a tidal radius, so for this profile, we will instead use the effective tidal radius, $r_{\rm te}$, as defined in Equation (2.3).
@@
With these definitions, Figure 4 shows the evolution of the scale radius, the virial radius, the tidal radius, and the two concentration parameters as the bound mass fraction decreases from left to right on the $x$ axis. For all three Energy Truncation models (NFWT, EinHighT and EinLowT) the scale radius decreases monotonically as the halo is tidally stripped, as expected [27]. The relative change in $r_{-2}$ is always slower than the relative mass loss rate, however, such that when the system has lost 90%  of its mass, $r_{-2}$ has only decreased to 20–50 per cent of its initial value. The behaviour for the evolution of $r_{-2}$ the H03 approximation is similar, although this model generally predicts a scale radius that is 10–50 per cent higher than the Energy Truncation model. The ST approximation predicts a constant $r_{-2}$ until the tidal radius is less than the scale radius (at approximately $0.5\,M_{\rm unit}$), at which point the evolution of the scale radius is similar to the Energy Truncation models.
@@
Figure 4: Evolution of the parameters used to characterize subhalo concentration as a function of mass fraction, for various profiles: (1) the scale radius $r_{-2}$, (2) the virial radius $r_{\rm vir}$; the radius enclosing an overdensity of $\bar{\rho}_{\rm vir}=0.2387\,\rho_{\rm unit}$ (3) the tidal radius (4) the concentration parameter $c=r_{\rm vir}/r_{-2}$, and (5) the effective concentration $c_{t}=r_{t}/r_{-2}$. All three Energy Truncation models evolve similarly to each other, with the EinLowT concentration being about 20 per cent higher than the NFWT, and EinHighT models. The concentration parameters calculated from two approximate models, H03 and ST, evolve quite differently.
@@
The tidal radius, $r_{t}$, also decreases with mass, however—unlike $r_{\rm vir}$—the value of $r_{t}$ varies widely between the models. For the three Energy Truncation models, $r_{t}$ is initially very large and decreases rapidly until the mass is approximately 10 per cent of $M_{\rm unit}$. At smaller masses, the three Energy Truncation models agree quite well. The H03 and ST approximations do not agree with Energy Truncation models and, instead, look similar to each other. Both of these approximations start with $r_{t}=r_{\rm vir}$ and then rapidly decrease. We emphasize that the tidal radius is defined differently for the H03 model, using the parameter $r_{te}$. This effective tidal radius is used in the H03 parameterization but does not have a clear physical meaning. We also note that the “virial" radius is typically larger than the tidal radius, except for the Energy Truncation models at high mass fractions; this means the virial radius is simply given by $r_{\rm vir}=(4\pi\bar{\rho}_{\rm vir}/3M)^{1/3}$.
@@
Taking the ratio of the two outer radii to the scale radii, we calculate corresponding concentration parameters. For the Energy Truncation models, $c=r_{r_{\rm vir}}/r_{-2}$, increases with decreased mass. We emphasize however that this apparent increase in $c$ comes from the definition of the virial radius, which tracks the original density of the system and ends up outside the actual tidal radius. Physically, the material still bound to the subhalo within its tidal radius is less concentrated — this is better captured by the second concentration parameter, $c_{t}$. We show how the second effective concentration parameter varies with mass loss in the bottom panel of Figure 4. Once the subhaloes have been stripped to about 20% of their original mass, the effective concentration is approximately constant at $c_{t}\approx 10$. Given the faster decline in its scale radius, the EinLowT profile has higher concentrations than the NFWT and EinLowT profiles. The ST and H03 models under-predict $c_{t}$ by an order of magnitude or more, though in the case of the H03, this is largely because the tidal radius definition is different.
@@
Overall, all profiles and mass loss models show certain broad trends. Stripped subhaloes have higher densities than field haloes of the same mass, but their bound material is more uniformly distributed than that of field haloes. Thus, concentrations relative to the original virial boundary generally increase, while concentrations relative to the tidal radius decrease. Nonetheless, there significant differences in concentration evolution are predicted simply based on differences in (1) the initial profile model and (2) the tidal stripping model. Finally, we note that the Energy Truncation model predicts that at low bound mass fractions, the scale radius and tidal radius will evolve at similar rates, resulting in a constant $c_{t}$ value.
@@ 4 Dark matter annihilation
One potential technique for determining the identity of dark matter particles is through dark matter annihilation. To place constraints on particle masses and interaction cross-sections, it is important to have accurate predictions of the distribution of dark matter within subhaloes, since the dark matter annihilation signals depend sensitively on substructure. Assuming the dark matter particles are Majorana WIMPs (weakly interacting massive particles) that can annihilate with one another, the rate at which dark matter annihilates is given by:
@@
where $\langle\sigma v\rangle$ is the velocity-averaged annihilation cross section, $m$ is the mass of the dark matter particle, $\rho$ is the density and $V$ is a volume [e.g. 80]. Thus, the annihilation rate is proportional to the quantity
$J\equiv\int\rho^{2}{\rm d}V\,\,\,,$ (4.2)
which is sometimes called the $J$-factor [^3] Here we follow the definition of [14]; more commonly the $J$-factor is expressed as a projected quantity., and has dimensions mass $\times$ density. This quantity is sensitive to the high-density inner region of the halo and does not vary with the choice of integration volume as long as the volume is large enough to enclose the region of high density.
@@
which is often called the boost factor. [^4] Not to be confused with the substructure boost factor, which is the luminosity of the signal divided by the luminosity from the smooth host halo [68]. Unlike the annihilation rate, the boost factor is sensitive to the integration volume. In the case of CDM haloes, it is common to use the spherical volume within the virial radius. The boost factor is useful in calculating the total annihilation signal from a system; a common approach is to draw subhalo properties such as infall redshift, mass, concentration, and/or orbital properties from random distributions, and then use the boost factor to calculate the contribution from each subhalo to the total signal [e.g. 4, 37, 77, 68, 40, 43, 44, 17, 76, 34].
@@
In terms of the boost factor, the annihilation rate can be expressed as:
$R(V)=\dfrac{\langle\sigma v\rangle}{2m^{2}}\bar{\rho}^{2}VB(V)\,\,\,,$ (4.4)
@@
where the first term, $\langle\sigma v\rangle/2m^{2}$, depends on the particle physics of the dark matter candidate, the second term $\bar{\rho}^{2}V$ depends on the mean density and volume (or equivalently the mean density and total mass) of the system, and $B(V)$ characterizes the inhomogeneity of the particle distribution within the volume $V$.
@@
The boost factor (top) and $J$-factor (bottom) calculated within a spherical volume of radius $r$ are shown in Fig 5 for each profile model and various degrees of mass loss. For the boost factor (which is a measurement of how inhomogeneous the mass distribution is), the parameter is roughly constant within the scale radius, $r_{-2}$, and then increases rapidly with radius (with $B\propto V$ outside the tidal radius). Inside $r_{-2}$, there is little evolution in the boost factor as mass is stripped, while outside this radius, the boost factor is higher for more stripped systems. Conversely, the $J$-factor increases with a radius for $r<r_{-2}$ and then is approximately constant at larger radii, but its value decreases as the system is stripped, especially in model H03.
@@
Figure 5: The boost factor (top) and $J$-factor (bottom) for tidally stripped subhalo models, as a function of the radius of the spherical volume within which the integral is calculated. The colour scale indicates the degree of stripping. The characteristic radii $r_{-2}$, $r_{\rm vir}$ and $r_{t}$ are labelled as circles, triangles, and crosses, respectively. The boost factor is approximately constant within the scale radius and then increases as $r^{3}$. Conversely, the $J$-factors increase rapidly within the scale radius and then are constant at large radii. This plot demonstrates how the boost factor and $J$-factor measurements depend on the choice of integration volume.
4.1 Boost Factor Evolution
@@
In the top row of Figure 6, we show how the boost factor evolves as a function of bound mass using different volumes. As expected from Figure 5, when calculated within the scale radius, the boost factor stays roughly constant. Within the virial radius— except the H03 approximation—the boost factor increases as mass is stripped. Finally, within the tidal radius, the boost factor decreases as the profile is stripped, with both the H03 and ST approximation having a lower boost factor by two orders of magnitude. This difference is largely because of the different volumes in which the boost factor is calculated. If we focus on the evolution of the boost factor relative to its initial value (second row), this difference disappears.
@@
Figure 6: Evolution of the boost factor calculated within different volumes. The top row shows the boost factor, the second panel shows the evolution from the initial value, and the bottom shows the ratio between the boost factor of the tidally-stripped profiles and the initial, untruncated profiles.
@@
It is also worth comparing predictions of boost factors to the values obtained using the untruncated profile. The bottom row of Figure 6 shows the ratio $B/B_{\rm un}$, where $B_{\rm un}$ is the initial un-truncated profile. We compare our findings to the constant value of 1.3, since [80] showed that the H03 model follows the relation $B\simeq 1.3\,B_{\rm un}$ when measured within $r_{te}$.
@@
When measured inside the scale radius, the boost factor is constant for all models and consistent with $B=1.3B_{\rm un}$. When measured within the virial radius, we find much higher values of $B/B_{\rm un}$. This result is because the virial radius is typically larger than the tidal radius in these systems. Since mass loss reduces the density in the outer parts of the stripped halo, this makes the density distribution of stripped systems more centrally concentrated and less homogenous, and thus it increases $B$.
@@
When measured within the tidal radius, we find the ST and H03 approximations agree with the $B\simeq 1.3\,B_{\rm un}$ prediction, as expected. The more accurate Energy Truncation stripping models predict a boost factor that is more than 10 times higher by the time the bound mass fraction is less than 10 per cent. We conclude that the boost factor is extremely sensitive to the details of mass loss within the tidal radius, and may have been underestimated in previous work. Thus, it may not provide the best way to calculate the annihilation signal

### v3

@@ 3 Evolution of the concentration parameter
Density profiles can be characterized by a concentration parameter, which reflects the distribution of mass within the system. For isolated haloes, the concentration parameter $c$ is traditionally defined in terms of the radius at which the logarithmic density profile has a slope of $-2$, i.e., $c=r_{\rm vir}/r_{-2}$ [35], where $r_{\rm vir}$ is the virial radius. As discussed in [39, 51], this definition does not capture deviations from an NFW profile. However, concentration is broadly correlated with formation epoch, early-forming haloes being more concentrated, though the details of the relationship are complicated [e.g. 52, 53, 54, 55, 56, 57, 58].
@@
In cosmological simulations, the virial radius is normally defined in terms of a (redshift-dependent) mean enclosed density. In idealized merger studies, the definition of the virial radius is somewhat arbitrary, but as in Section 2.3, we assume the satellite haloes begin merging with a concentration of $c=10$. This corresponds to an enclosed density of $\bar{\rho}_{\rm vir}=0.2387\,\rho_{\rm unit}$. As the satellite is stripped, we solve for $r_{\rm vir}$ as the radius where $\bar{\rho}(r_{\rm vir})=\bar{\rho}_{\rm vir}$. The scale radius, $r_{-2}$, is calculated directly by numerically differentiating and then solving ${\rm d}\log_{10}\rho/{\rm d}\log_{10}r=-2$. For the ST profile, once the tidal radius (defined below) is within the original scale radius, we use $r_{-2}=r_{t}$.
@@
As we will show below, this first definition of concentration generally increases as mass is lost since the scale radius shrinks faster than $r_{\rm vir}$. While this definition is useful for tracking how the bound remnant of a subhalo increases in mean density as it loses its lower-density outer material, $r_{\rm vir}$ has no clear physical significance once a halo has become a subhalo since its density no longer tracks the original cosmological mean value. Thus, in addition to the classical concentration, $c=r_{\rm vir}/r_{-2}$, for stripped systems we will consider an “effective" concentration parameter, $c_{t}=r_{t}/r_{-2}$. As argued in [59, 2], this definition may be more natural as it better reflects the higher mean density of subhaloes and the lower concentration the subhalo had when it was originally accreted. In general, we define the tidal radius $r_{t}$ to be the radius at which the density profile is zero. With this definition, the H03 profile does not have a tidal radius, so for this profile, we will instead use the effective tidal radius, $r_{\rm te}$, as defined in Equation (2.3).
@@
With these definitions, Figure 4 shows the evolution of the scale radius, the virial radius, the tidal radius, and the two concentration parameters as the bound mass fraction decreases from left to right on the $x$ axis. For all three Energy Truncation models (NFWT, EinCoreT and EinCuspT) the scale radius decreases monotonically as the halo is tidally stripped, as expected [25]. The relative change in $r_{-2}$ is always slower than the relative mass loss rate, however, such that when the system has lost 90%  of its mass, $r_{-2}$ has only decreased to 20–50 per cent of its initial value. The behaviour for the evolution of $r_{-2}$ in the H03 approximation is similar, although this model generally predicts a scale radius that is 10–50 per cent higher than the Energy Truncation model. The ST approximation predicts a constant $r_{-2}$ until the tidal radius is less than the scale radius (at approximately $0.5\,M_{\rm unit}$), at which point it rapidly decreases.
@@
Figure 4: Evolution of the parameters used to characterize subhalo concentration as a function of bound mass fraction, for various profiles: (1) the scale radius $r_{-2}$, (2) the virial radius $r_{\rm vir}$; the radius enclosing an overdensity of $\bar{\rho}_{\rm vir}=0.2387\,\rho_{\rm unit}$, (3) the tidal radius, (4) the concentration parameter $c=r_{\rm vir}/r_{-2}$, and (5) the effective concentration $c_{t}=r_{t}/r_{-2}$. All three Energy Truncation models evolve similarly to each other, with the EinCuspT concentration being about 20 per cent higher than the NFWT, and EinCoreT models. The concentration parameters calculated from the other two models, H03 and ST, evolve quite differently.
@@
The tidal radius, $r_{t}$, also decreases with mass, however—unlike $r_{\rm vir}$—the value of $r_{t}$ varies widely between the models. For the three Energy Truncation models, $r_{t}$ is initially very large and decreases rapidly until the mass is approximately 10 per cent of $M_{\rm unit}$. At smaller masses, the three Energy Truncation models agree quite well. The H03 and ST approximations do not agree with Energy Truncation models and, instead, look similar to each other. Both of these approximations start with $r_{t}=r_{\rm vir}$ and then rapidly decrease. We emphasize that the tidal radius is defined differently for the H03 model, using the parameter $r_{te}$. This effective tidal radius is used in the H03 parameterization but does not have a clear physical meaning. We also note that the “virial” radius is typically larger than the tidal radius, except for the Energy Truncation models at high mass fractions; this means the virial radius is simply given by $r_{\rm vir}=(4\pi\bar{\rho}_{\rm vir}/3M)^{1/3}$.
@@
Taking the ratio of the two outer radii to the scale radii, we calculate the corresponding concentration parameters. For the Energy Truncation models, $c=r_{r_{\rm vir}}/r_{-2}$, increases with decreased mass. We emphasize, however, that this apparent increase in $c$ comes from the definition of the virial radius, which tracks the original density of the system and will extend beyond the tidal radius. Physically, the material still bound to the subhalo within its tidal radius is less concentrated—this is better captured by the second concentration parameter, $c_{t}$. We show how the second effective concentration parameter varies with mass loss in the bottom panel of Figure 4. Once the subhaloes have been stripped to about 20% of their original mass, the effective concentration is approximately constant at $c_{t}\approx 10$. Given the faster decline in its scale radius, the EinCuspT profile has higher concentrations than the NFWT and EinCuspT profiles. The ST and H03 models predict a lower $c_{t}$ by an order of magnitude or more compared to the Energy Truncation model; though in the case of the H03, this is largely because the tidal radius definition is different.
@@
Overall, all profiles and mass loss models show certain broad trends. If we take the mass of field haloes to be their virial mass within $r_{\rm vir}$, and the mass of stripped subhaloes to be the bound mass within the tidal radius $r_{t}$, then stripped subhaloes have higher mean densities than field haloes of the same mass, but their bound material is more uniformly distributed than that of field haloes. Thus, concentrations relative to the original virial boundary generally increase, while concentrations relative to the tidal radius decrease. Nonetheless, significant differences in concentration evolution are predicted simply based on differences in (1) the initial profile model and (2) the tidal stripping model. Finally, we note that the Energy Truncation model predicts that at low bound mass fractions, the scale radius and tidal radius will evolve at similar rates, resulting in a constant $c_{t}$ value.
@@ 4 Dark matter annihilation
One potential technique for determining the identity of the dark matter particle is through dark matter annihilation. To place constraints on particle masses and interaction cross-sections, it is important to have accurate predictions of the distribution of dark matter within subhaloes since the dark matter annihilation signal depends sensitively on substructure. A common approach to modelling the contribution of substructure to annihilation signals is to draw subhalo properties such as infall redshift, mass, concentration, and/or orbital properties from random distributions or known scaling relations, and then to integrate the contribution from each subhalo to the total signal [e.g. 59, 60, 61, 2, 62, 63, 64, 65, 66, 67].
Assuming the dark matter particles are Majorana WIMPs (weakly interacting massive particles) that can annihilate with one another, the rate at which dark matter annihilates is given by:
@@
where $\langle\sigma v\rangle$ is the velocity-averaged annihilation cross section, $m$ is the mass of the dark matter particle, $\rho$ is the density and $V$ is a volume [e.g. 68]. Thus, the annihilation rate is proportional to the quantity
$J(V)\equiv\int_{V}\rho^{2}{\rm d}V^{\prime}\,\,\,,$ (4.2)
which is sometimes called the $J$-factor444 Though the $J$-factor is traditionally expressed as the flux along the line-of-sight [69], here we follow the definition of [70], where the $J$-factor reflects the absolute annihilation luminosity. , and has dimensions mass×$\times$×density. This quantity is sensitive to the high-density inner region of the halo and does not depend strongly on the choice of integration volume, as long as that volume is large enough to enclose the region of high density.
@@
where $\bar{\rho}$ is the mean density within the volume $V$. We will refer to this as the “flux multiplier”, as in [68]. [^5] This quantity has also been called the halo boost factor [2]. Unlike the annihilation rate, this quantity is sensitive to the integration volume. In the case of CDM haloes, it is common to use the spherical volume within the virial radius.
@@
In terms of the flux multiplier, the annihilation rate for a single halo can be expressed as:
$R(V)=\dfrac{\langle\sigma v\rangle}{2m^{2}}\bar{\rho}^{2}Vf_{M}(V)\,\,\,,$ (4.4)
@@
where the first term, $\langle\sigma v\rangle/2m^{2}$, depends on the particle physics of the dark matter candidate, the second term $\bar{\rho}^{2}V$ depends on the mean density and volume (or equivalently the size and total mass) of the system, and $f_{M}(V)$ characterizes the mass profile of the halo within the volume $V$.
@@
The flux multiplier ($f_{M}$; top) and $J$-factor (J; bottom) calculated within a spherical volume of radius $r$ are shown in Fig 5 for each profile model and various degrees of mass loss. For the flux multiplier (which is a measurement of how inhomogeneous the mass distribution is), the parameter is roughly constant within the scale radius, $r_{-2}$, and then increases rapidly with radius (with $f_{M}\propto V$ outside the tidal radius). Inside $r_{-2}$, there is little evolution in the flux multiplier as mass is stripped, while outside this radius, the flux multiplier is higher for more stripped systems. Conversely, the $J$-factor increases with a radius for $r<r_{-2}$ and then is approximately constant at larger radii, but its value decreases as the system is stripped, especially in model H03.
@@
Figure 5: The flux multiplier, $f_{M}(r)$ (top) and $J$-factor, $J(r)$ (bottom) for tidally stripped subhalo models, as a function of the radius of the spherical volume within which the integral is calculated. The colour scale indicates the degree of stripping. The characteristic radii $r_{-2}$, $r_{\rm vir}$ and $r_{t}$ are labelled as circles, triangles, and crosses, respectively. The flux multiplier is approximately constant within the scale radius and then increases as $r^{3}$. Conversely, the $J$-factors increase rapidly within the scale radius and then are constant at large radii. This plot demonstrates how the flux multiplier and $J$-factor measurements depend on the choice of integration volume.
4.1 Flux multiplier evolution
@@
In the top row of Figure 6, we show how the flux multiplier evolves as a function of bound mass using different volumes. As expected from Figure 5, when calculated within the scale radius, the flux multiplier stays roughly constant. Within the virial radius—except the H03 approximation—the boost factor increases as mass is stripped. Finally, within the tidal radius, the flux multiplier decreases as the profile is stripped, with both the H03 and ST approximation having a lower flux multiplier than the Energy Truncation Model by two orders of magnitude. This difference is largely because of the different volumes in which the flux multiplier is calculated. If we focus on the evolution of the flux multiplier relative to its initial value (second row), this difference disappears.
@@
Figure 6: Evolution of the flux multiplier calculated within different volumes. The top row shows the flux multiplier , the second panel shows the evolution from the initial value, and the bottom shows the ratio between the flux multiplier of the tidally-stripped profiles and the initial, untruncated profiles.
@@
It is also worth comparing predictions of flux multipliers to the values obtained using the untruncated profile. The bottom row of Figure 6 shows the ratio $f_{M}/f_{M,{\rm un}}$, where $f_{M,{\rm un}}$ is the initial un-truncated profile. We compare our findings to the constant value of 1.3, since [68] showed that the H03 model follows the relation $f_{M}\simeq 1.3\,f_{M,{\rm un}}$ when measured within $r_{te}$.
@@
When measured inside the scale radius, the flux multiplier is constant for all models and consistent with $f_{M}=1.3f_{M,{\rm un}}$. When measured within the virial radius, we find much higher values of $f_{M}/f_{M,{\rm un}}$. This result is because the virial radius is typically larger than the tidal radius in these systems. Since mass loss reduces the density in the outer parts of the stripped halo, this makes the density distribution of stripped systems more centrally concentrated and less homogenous, and thus it increases $f_{M}$.
@@
When measured within the tidal radius, we find the ST and H03 approximations agree with the $f_{M}\simeq 1.3\,f_{M,{\rm un}}$ prediction, as expected. The more accurate Energy Truncation stripping models predict a flux multiplier that is more than 10 times higher by the time the bound mass fraction is less than 10 per cent. We conclude that the flux multiplier is extremely sensitive to the details of mass loss within the tidal radius, and may have been underestimated in previous work.